Glacial acetic acid is available as a 90% weight by weight, specific gravity is 1.255. what volume of the acid will be needed to prepare a 1000mill of 10millimolar solution assuming a formula weight of 80
1
Expert's answer
2021-08-31T01:32:02-0400

CM = n/V

0.01 = n / 1 L

n = 0.01 x 1 = 0.01 mol

n = m / M

m = n x M = 0.01 x 80 = 0.8 g

Mass Glacial acetic acid of 90% purity (0.8 x 100) / 90 = 0.89 g

density = m/V

V = m / density = 0.89 / 1.255 = 0.71 ml